# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

### v3

@@ 3.5 C5: Whale Re-Identification
Figure 11: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "white underside of humpback whale fluke" where the fluke was clearly visible and of sufficient quality for pattern matching. (Bottom) Matching individuals to the HappyWhale database and analysis. A deep learning Re-ID model retrieves candidate matches in the HappyWhale database, which we manually verify. We then map repeated sightings of confirmed individuals, demonstrating how INQUIRE-Search can support analyses of movement patterns and habitat use.
Of the 153 candidate images retrieved and verified through INQUIRE-Search, 57 were successfully matched to individuals in the public HappyWhale dataset. Despite inspecting only the top 200 ranked search results, these matches identified 34 unique humpback whales, including multiple sightings of a few individuals, as shown in Fig 11. These additional sightings extend spatiotemporal coverage of known individuals, demonstrating that INQUIRE-Search can uncover valuable, previously untapped data embedded within large community science repositories.
@@
Fig 11 also visualizes matched iNaturalist and HappyWhale observations, highlighting how this linking expands our ability to track individuals across datasets. Since iNaturalist observations include geolocation metadata, cross-referencing matches provides new spatial information for individuals that were previously represented only in curated catalogs. This capability enables researchers to map additional movement paths, identify unusual sightings, and potentially detect shifts in habitat use when opportunistic images align with structured monitoring efforts.
Findings
While small in scale, this case study demonstrates that images can be surfaced by INQUIRE-Search, then linked to known individuals using re-ID, augmenting existing datasets and providing valuable new information for population monitoring. INQUIRE shows strong potential for retrieving high-quality, correctly posed images that capture identifiable features, which enables the integration of valuable data from previously disconnected sources. However, its broader utility is constrained by the identification step, which still depends on a pre-existing labeled dataset and a trained model for matching. Because INQUIRE cannot assign individual identities on its own, its use is more limited when building datasets from scratch or incorporating previously unseen individuals. In addition, retrieval performance reflects the spatial and behavioral biases of community-sourced photographs, meaning the tool will preferentially surface observations from well-sampled areas and charismatic encounters rather than providing an unbiased representation of whale distribution or behavior.
@@ 4 Discussion
Across our five case studies, INQUIRE-Search enabled rapid, expert-guided discovery by turning community-sourced images into analyzable ecological data. The system consistently recovered information that would have required substantial fieldwork, helping to resolve major temporal, spatial, and observational gaps.
4.1 Rethinking the Scientific Process in the Age of Image Search
@@
With the interactive, iterative approach to data discovery enabled by INQUIRE-Search, scientists remain the drivers of inquiry, interpretation, and verification. Incorporating such a tool with an ecological research workflow requires rethinking experimental design. The following discussion highlights parameters to "tune" in this process, from identifying suitable questions and crafting precise prompts, to allocating verification effort and accounting for uncertainty.
4.1.1 Which questions are well-posed for INQUIRE-Search?
@@
INQUIRE-Search is fundamentally limited by the images that exist in the underlying repository. Questions targeting extremely rare events, under-observed species, or obscure behaviors may yield insufficient data for robust analysis. For example, in the avian diet case study, species with few iNaturalist records produced too few relevant images of feeding behaviors to support quantitative comparisons across time. Conversely, when data are overly-abundant, as in the wildlife mortality study, search still improves efficiency but does not eliminate the need for substantial verification. Useful queries must also be unambiguously visually verifiable: prompts tied to explicit, observable features (“milkweed with seed pods”) are more reliable than abstract states (“tree under stress”). Well-posed questions therefore target phenomena that are (1) sufficiently represented in the data pool, (2) not so broad or ubiquitous that they produce large volumes results, and (3) visually verifiable.
4.1.2 How to refine a question: iterating on prompts
@@
Search-based discovery depends on clear, specific prompts. Precise language reduces ambiguity in the embedding space and improves the relevance of returned images. Specificity also increases the verifiability of the outputs: "robin pulling a worm from the soil" will yield cleaner results than "bird foraging. Iteration is often necessary, where prompts can be refined by adding behavioral, contextual, or anatomical cues, while overly narrow prompts can be relaxed when returns are sparse.
4.1.3 Managing Effort, Ranking, and Verification
@@
This workflow introduces a new design choice: in addition to formulating effective queries, researchers must also decide how far down the ranked search results to inspect. In this way, "effort" becomes tunable. Reviewing only the top-ranked subset often recovers the most relevant observations with minimal redundancy, as shown in the dietary analysis case study, where experts verified only first 200 images recovered to obtain a sufficient quantity of data.
4.1.4 Identifying Uncertainty and Bias with INQUIRE-Search
@@
INQUIRE-Search inherits the biases of community-science platforms. Observer preferences skew data toward charismatic, accessible, or unusual events, while cryptic interactions and remote habitats remain under-represented [dimson2023observer]. Although methods for observer-effort corrections to quantify and mitigate these biases exist [robinson2018correcting, guilbaultexplorers, mondain2024adaptive, padilla2024assessing], their effectiveness at a global scale remains limited because they require structured sampling, repeated surveys, or explicit effort measurements that community science data often lack.
@@
Model-driven biases add a second layer of uncertainty. VLMs also struggle with prompts requiring complex relational or compositional understanding [alhamoud2025visionlanguagemodelsunderstandnegation, thrush2022winogroundprobingvisionlanguage]. Moreover, recent work shows that VLMs are both miscalibrated (where the expressed confidence does not match true accuracy) [guo2017calibration] and biased (the model defaults to memorized training-set associations instead of analyzing visual evidence) [vo2025vision]. Adapting existing statistical tools for bias estimation and correction [narduzzi2014inverse, geifman2017selective, karimi2023quantifying] could provide a foundation for quantifying uncertainty in discovery-driven search.
@@ 4.2 Expanding Beyond iNaturalist
Although our case studies drew on iNaturalist for its scale and ecological breadth, the open-source design of the INQUIRE-Search codebase makes it adaptable to any large image collection. Early qualitative experiments suggest that the system has potential for other ecological image repositories, such as camera trap datasets. Because the system is open-source and resource-efficient, it has the potential to be deployed widely, enabling scientists, NGOs, and local communities to harness image search for ecological monitoring and hypothesis generation on their own databases.
@@ 5 Conclusion
Growing ecological image databases remain an underutilized scientific resource. We demonstrate that INQUIRE-Search enables iterative, natural language–driven discovery and verification of diverse ecological signals, supporting scalable analysis of phenomena such as bird diet and plant phenology. This approach suggests a new paradigm in ecological research, in which hypothesis generation, testing, and data curation are integrated within an interactive search system over existing data. By transforming community-sourced imagery from incidental observations into systematic ecological evidence, INQUIRE-Search broadens access to data-driven discovery while raising important questions about rigor, reproducibility, and inference. When combined with careful experimental design, it offers a scalable pathway to accelerate ecological insight in a changing world
Supplementary Information
5.1 System Architecture and Design Motivation
@@
INQUIRE-Search is a tool designed to enable scientists to quickly and easily discover data from within a large ecological database such as iNaturalist. This challenge of surfacing data of interest from a large pool is commonly referred to as "image retrieval" in the AI community, and INQUIRE-Search specifically addresses two fundamental image retrieval challenges to ensure the system is high-performing in terms of both quality and speed. First, INQUIRE-Search enables scientists to search for data using open-ended text, which significantly expands the searchable possibilities beyond existing metadata within the database. Second, INQUIRE-Search is engineered to power this search over very large databases efficiently, without requiring significant computational resources. This efficiency is a key feature of the design, letting scientists iterate on their searches with the system in real-time to discover their data of interest. The system architecture of INQUIRE-Search reflects both of these priorities, combining state-of-the-art vision-language models with efficient indexing and memory management techniques to deliver responsive searches across hundreds of millions of images on modest hardware (4 vCPUs with 32GB RAM, roughly the capacity of a single laptop).
5.2 Search Index
@@
Vector databases, which use a specialized indexing system that allows fast search over millions of high-dimensional vectors, [decastro2018detecting, singla2021experimental, fang2017topology] are specifically designed to facilitate efficient similarity-based search. We utilize FAISS [douze2025faiss] to create an approximate nearest neighbor index that is small and fast to query. Our FAISS index is generated and tuned using the Autofaiss library, a widely-used library for efficient nearest-neighbor search [webster2023duplication]. The final index is memory-mapped (stored on disk but accessed as if it were in memory, enabling fast search with limited RAM), uses 36GB of storage, and enables sub-500ms search times while maintaining high retrieval accuracy.
@@
5.3 Image and Text Embeddings
The core search capability behind INQUIRE-Search is powered by such a vision-language model. Not all vision-language models are equally capable of ranking ecological data, particularly for scientific queries. We leveraged our previous work developing the INQUIRE-Benchmark, which specifically evaluates models for ecological image retrieval tasks, to select SigLIP-So400m-384-14 [zhai2023sigmoid] as our embedding model, due to both its strong performance on the benchmark and its reasonable tradeoff between speed and accuracy; other models like ViT-H-14 [radford2021learning] are slightly more accurate at retrieving relevant data but are significantly larger, and thus slower. We processed 300M images sourced from iNaturalist through the visual encoder component of the selected SigLIP model to build the backbone of the INQUIRE-Search tool.
Figure 12: INQUIRE-Search user interface. INQUIRE-Search provides an user-friendly and intuitive interface that allows ecologists to query hundreds of millions of natural-world images. This example illustrates a query for "American robin eating," with species and temporal filters applied, enabling rapid retrieval, inspection, and verification of relevant observations.
5.4 Search Interface
INQUIRE-Search provides a web-based interface (Fig 12) that surfaces the search index and filters to scientists. The interface includes a query field for natural language queries (e.g., "California condor with a green ‘26’ tag on its wing") and controls for taxonomic, temporal, or geospatial filters. The interface displays a grid of thumbnails of the images returned from the user-provided query, ordered by relevance to the query. The images are clickable, allowing the user to mark images for further analysis. An expanded view of a selected image shows a full-resolution image with complete metadata.
5.5 Data Export
@@
5.6 Data Analysis
Once verified, these retrieved observations form the foundation for a wide range of statistical analyses, demonstrated in the following section. This workflow allows experts to effectively "design" their search strategy and specify what ecological signals to target, while identifying or mitigating relevant sources of uncertainty and bias.
5.7 Retrieval and Filtering Results
@@
American Robin "<species> with <diet type> in its mouth" Summer 500 per type 338 67.6%
@@
Young coniferous trees "young coniferous trees in burned forest" Geo: High Park Fire (40.57-40.75 N, 105.18-105.54 W); Date: Post-2012 100 78 78.0%
Young deciduous trees "young deciduous trees in burned forest" 112 45 40.2%
@@
Boston (Urban) "dead bird" Geo: Bounding box around Boston, MA; No Taxon filter 1000 543 54.3%
Pioneer Valley (Rural) "dead bird" Geo: Bounding box around Pioneer Valley/Amherst, MA; No Taxon filter 360 79 21.9%
@@
Emergence "Milkweed germinating or emerging" Taxon: Asclepias syriaca; Geo: S. Quebec (45.03 to 46.54 N, -74.68 to -71.66 W); No temporal filter 200 45 22.5%
Flowering "Milkweed flowering" 200 169 84.5%
Seeding "Milkweed producing seeds…" 200 161 80.5%
Senescence "Milkweed dying or withering…" 200 52 26.0%
@@
Humpback whale flukes "white underside of humpback whale fluke" Taxon: Humpback Whale (Megaptera novaeangliae); No Geo/Temp filters 200 153 76.5%

### v5

@@ 3.5 CS5: Whale re-identification
Figure 11: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for “white underside of humpback whale fluke” where the fluke was clearly visible and of sufficient quality for pattern matching. (Bottom) Matching individuals to the HappyWhale database and analysis. A deep learning Re-ID model retrieves candidate matches in the HappyWhale database, which we manually verify. We then map repeated sightings of confirmed individuals, demonstrating how INQUIRE-Search can support analyses of movement patterns and habitat use.
@@
Using the verified INQUIRE-Search images, we matched 57 observations to the HappyWhale dataset, corresponding to 34 unique humpback whales (Figure 11, top). Several individual appeared multiple times, extending their spatiotemporal coverage. Figure 11 (bottom) visualizes matched iNaturalist and HappyWhale observations, highlighting how this linking expands our ability to track individuals across datasets. Since iNaturalist observations include geolocation metadata, cross-referencing matches provides new spatial information for individuals that were previously represented only in curated catalogs. This capability enables the identification of additional movement paths and unusual sightings.
INQUIRE-Search enables efficient discovery of correctly posed, identifiable fluke images from unstructured community-science imagery, facilitating scalable integration with existing photo-ID datasets. However, downstream identification remains dependent on pre-existing labeled catalogs and trained re-identification models. Also, retrieval reflects spatial and behavioral biases in community-sourced photography, favoring well-sampled locations and charismatic encounters over unbiased representations of whale distribution or behavior.
@@ 4 Discussion
INQUIRE-Search enables rapid, expert-guided discovery by efficiently surfacing analyzable ecological data from community-sourced images. The case studies presented in this work serve as concrete examples of how this new capability can be integrated into scientific workflows, illustrating how questions are posed, prompts refined, verification effort allocated, and downstream analyses conducted. Across diverse applications, we demonstrate that the system consistently recovers information that would otherwise require substantial manual effort or targeted fieldwork, helping to resolve temporal, spatial, and observational gaps.
4.1 INQUIRE-Search in practice: design considerations and limitations
@@
INQUIRE-Search reframes the early stages of the scientific process by transforming large image archives from passive repositories into interactive substrates for reasoning. Scientists remain the drivers of inquiry, interpretation, and verification, but the bottleneck shifts from data collection to validation. Incorporating such a tool with an ecological research workflow requires rethinking experimental design. The following discussion highlights parameters to “tune” in this process, from identifying suitable questions and crafting precise prompts, to allocating verification effort and accounting for uncertainty.
@@
Which questions are well-posed for INQUIRE-Search? INQUIRE-Search is fundamentally limited by the images that exist in the underlying repository. Questions targeting rare events, under-observed species, or obscure behaviors may yield insufficient data for robust analysis. For example, in the avian diet case study, species with few iNaturalist records produced too few relevant images of feeding behaviors to support quantitative comparisons across time. Conversely, when data are overly abundant, as in the wildlife mortality study, search still improves efficiency but does not eliminate the need for substantial verification. Useful queries must also be unambiguously visually verifiable: prompts tied to explicit, observable features (“milkweed with seed pods”) are more reliable than abstract states (“tree under stress”). Well-posed questions therefore target phenomena that are (1) sufficiently represented in the data pool, (2) not so broad that they produce large volumes results, and (3) visually verifiable.
@@
How to refine a question: iterating on prompts. Search-based discovery depends on clear and specific prompts. Precise language reduces ambiguity in the embedding space and improves the relevance of returned images. Specificity also increases the verifiability of the outputs: “robin pulling a worm from the soil” will yield cleaner results than “bird foraging.” Iteration is often necessary, where prompts can be refined by adding behavioral, contextual, or anatomical cues, while overly narrow prompts can be relaxed when returns are sparse.
@@
Managing effort, ranking, and verification. This workflow introduces a new design choice: in addition to formulating effective queries, researchers must also decide how far down the ranked search results to inspect. In this way, “effort” becomes tunable. Reviewing only the top-ranked subset often recovers the most relevant observations with minimal redundancy, as shown in the dietary analysis case study, where experts verified only the first 200 images recovered to obtain a sufficient quantity of data.
@@
Identifying uncertainty and bias with INQUIRE-Search. INQUIRE-Search inherits the biases of community-science platforms. Observer preferences skew data toward charismatic, accessible, or unusual events, while cryptic interactions and remote habitats remain under-represented (Dimson and Gillespie, 2023). Although methods for observer-effort corrections to quantify and mitigate these biases exist (Robinson et al., 2018; Guilbault et al., 2025; Mondain-Monval et al., 2024; Padilla-Pozo et al., 2024), their effectiveness at a global scale remains limited because they require structured sampling, repeated surveys, or explicit effort measurements that community science data often lack.
@@
Model-driven biases add a second layer of uncertainty. VLMs also struggle with prompts requiring complex relational or compositional understanding (Alhamoud et al., 2025; Thrush et al., 2022). Moreover, recent work shows that VLMs are both miscalibrated (where the expressed confidence does not match true accuracy) (Guo et al., 2017) and biased (the model defaults to memorized training-set associations instead of analyzing visual evidence) (Vo et al., 2025). Adapting existing statistical tools for bias estimation and correction (Narduzzi et al., 2014; Geifman and El-Yaniv, 2017; Karimi and Samavi, 2023) could provide a foundation for quantifying uncertainty in discovery-driven search.
INQUIRE-Search supports dataset aggregation rather than end-to-end statistical inference. It retrieves candidate observations from large, unstructured image repositories via ranked search, leaving questions of representativeness, validation, and analysis to the researcher. Because both the data and retrieval process are opportunistic, results reflect community, model, and prompt biases and should not be treated as statistically representative samples. With experts in the loop, search outputs are treated as hypotheses rather than ground truth, placing responsibility on researchers to communicate uncertainty and limitations in downstream analyses.
4.2 Computation cost
INQUIRE-Search is built on embedding and retrieval rather than repeated model inference, substantially reducing computational and energy costs once the system is deployed. Images in the databases are embedded by the AI model once using a pretrained VLM and stored in an index, and each search requires only a single embedding of the user’s text query (single model inference call) followed by fast similarity matching. In contrast, other AI models such as generative models often reprocess all images and text in the database for each query, requiring repeated, compute-intensive inference that scales poorly with dataset size. The marginal energy cost of interactive discovery with INQUIRE-Search is low, making it a practical approach to large-scale search and retrieval.
@@ 4.3 Expanding beyond iNaturalist
Although our case studies drew on iNaturalist for its scale and ecological breadth, the open-source design of the INQUIRE-Search codebase makes it adaptable to any large image collection. Early qualitative experiments suggest that the system has potential for other ecological image repositories, such as camera trap datasets. The system’s novelty lies in the mechanism of discovery rather than the data source, enabling scientists, NGOs, and local communities to deploy the same open-source and resource-efficient workflow to harness image search for ecological monitoring and hypothesis generation on their own databases.
@@ 5 Conclusion
Growing ecological image databases remain an underutilized scientific resource. INQUIRE-Search, our proposed methodology and workflow for efficient and open-ended data discovery and verification via text-based search, enables scalable analysis of diverse ecological phenomena, from bird diet to plant phenology. This approach introduces a new paradigm in ecological research, in which hypothesis generation, experiment design, and data “collection” are integrated within an interactive search system over already-collected datasets. By enabling scientists to efficiently extract ecological evidence from community-sourced imagery, we can maximize the value of existing datasets. This can serve as an accessible and low-resource starting point for hypothesis testing, and will in turn enable the community to prioritize resource intensive data collection campaigns to fill remaining gaps. However, any new experimental methodology requires the development of best practices to ensure the resulting analyses are rigorous, reproducible, and unbiased. Through our case studies, we introduce initial considerations and discussion around careful experimental design with such a system. Our analyses show the diversity of potential applications of this approach and demonstrate the incredible efficiency at which experts can convert hypotheses to inference: this offers a scalable pathway to accelerate ecological insight in a rapidly changing world. By reducing discovery time across vast image databases, the system enables faster hypothesis generation and prioritization of follow-up studies. While search-based discovery does not replace systematic surveys, it can complement them by revealing patterns, gaps, and emerging signals that merit targeted investigation.
Glossary
Cosine Similarity: A metric measuring angular similarity between vectors, commonly used to rank semantic relevance.
Embedding Vector: A numerical representation of an image or text query used for similarity search and other downstream tasks.
Ground Truth: A reference dataset or expert-validated information used for comparison or validation.
Individual Re-Identification (Re-ID): The process of recognizing the same individual organism across multiple observations.
Joint Embedding Space: A high-dimensional vector space in which semantically similar images and text descriptions are located near one another.
Metadata Filtering: Restricting searches using structured fields such as taxonomy, date, or location.
Model Bias: Systematic errors introduced by training data or model assumptions that affect retrieval results.
Natural-Language Query: A free-text query written in plain language rather than predefined labels or categories.
Open-Set Retrieval: A retrieval setting where queries are not restricted to predefined classes or labels.
Opportunistic Sampling: Data collection driven by observer behavior rather than structured survey design.
Phenology: The timing of recurring biological events (e.g., flowering, senescence) across seasonal cycles.
Sampling Bias: Non-random representation of locations, species, or behaviors in collected data.
Secondary Data: Ecologically relevant information captured in images beyond species presence, such as behavior, interactions, or habitat context.
Vector Similarity Search: A retrieval method that ranks items based on distance (e.g., cosine similarity) between embedding vectors.
Verification Effort: The amount of expert review required to confirm retrieved images as informative.
Vision–Language Model (VLM): A machine learning model that embeds images and text into a shared semantic space, enabling cross-modal search and comparison.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author(s) used GPT-4o in order to catch grammatical errors and other inconsistencies in the writing. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Acknowledgements
We appreciate all the global contributors to the iNaturalist platform for their collection of species observations that make up the backbone of this work. We also want to thank the iNaturalist team and the authors of the INQUIRE benchmark, particularly Scott Laurie, Alex Shepard, Grant Van Horn, Omiros Pantazis, Gabriel Brostow, and Kate Jones. This work was supported in part by a Schmidt Science’s AI2050 Early Career Fellowship, NSF CAREER Grant (Award No. 2441060), an NSF Graduate Research Fellowship (Award No. DGE-2146755), the NSF and NSERC AI and Biodiversity Change Global Center (NSF Award No. 2330423 and NSERC Award No. 585136), and the MIT Generative AI Consortium. Oisin Mac Aodha was in part supported by a Royal Society Research Grant.
Appendix A: INQUIRE-Search system architecture and design motivation
@@ Acknowledgements
INQUIRE-Search is a tool designed to enable scientists to quickly and easily discover data from within a large ecological database such as iNaturalist. This challenge of surfacing data of interest from a large pool is commonly referred to as “image retrieval” in the AI community, and INQUIRE-Search specifically addresses two fundamental image retrieval challenges to ensure the system is high-performing in terms of both quality and speed. First, INQUIRE-Search enables scientists to search for data using open-ended text, which significantly expands the searchable possibilities beyond existing metadata within the database. Second, INQUIRE-Search is engineered to power this search over very large databases efficiently, without requiring significant computational resources. This efficiency is a key feature of the design, letting scientists iterate on their searches with the system in real-time to discover their data of interest. The system architecture of INQUIRE-Search reflects both of these priorities, combining state-of-the-art vision-language models with efficient indexing and memory management techniques to deliver responsive searches across hundreds of millions of images on modest hardware (4 vCPUs with 32GB RAM, roughly the capacity of a single laptop). A visualization of the tool is provided in Figure 12.
Figure 12: INQUIRE-Search provides an user-friendly and intuitive interface that allows ecologists to query hundreds of millions of natural-world images. This example illustrates a query for “American robin eating,” with species and temporal filters applied, enabling rapid retrieval, inspection, and verification of relevant observations.
Search index
@@
Vector databases, which use a specialized indexing system that allows fast search over millions of high-dimensional vectors, (DeCastro-García et al., 2018; Singla et al., 2021; Fang, 2017) are specifically designed to facilitate efficient similarity-based search. We utilize FAISS (Douze et al., 2025) to create an approximate nearest neighbor index that is small and fast to query. Our FAISS index is generated and tuned using the Autofaiss library, a widely-used library for efficient nearest-neighbor search (Webster et al., 2023). The final index is memory-mapped (stored on disk but accessed as if it were in memory, enabling fast search with limited RAM), uses 36GB of storage, and enables sub-500ms search times while maintaining high retrieval accuracy.
@@
Image and text embeddings
The core search capability behind INQUIRE-Search is powered by such a vision-language model. Not all vision-language models are equally capable of ranking ecological data, particularly for scientific queries. We leveraged our previous work developing the INQUIRE-Benchmark, which specifically evaluates models for ecological image retrieval tasks, to select SigLIP-So400m-384-14 (Zhai et al., 2023) as our embedding model, due to both its strong performance on the benchmark and its reasonable tradeoff between speed and accuracy; other models like ViT-H-14 (Radford et al., 2021) are slightly more accurate at retrieving relevant data but are significantly larger, and thus slower. We processed 300M images sourced from iNaturalist through the visual encoder component of the selected SigLIP model to build the backbone of the INQUIRE-Search tool.
Search interface
INQUIRE-Search provides a web-based interface (Figure 12) that surfaces the search index and filters to scientists. The interface includes a query field for natural language queries (e.g., “California condor with a green ‘26’ tag on its wing”) and controls for taxonomic, temporal, or geospatial filters. The interface displays a grid of thumbnails of the images returned from the user-provided query, ordered by relevance to the query. The images are clickable, allowing the user to mark images for further analysis. An expanded view of a selected image shows a full-resolution image with complete metadata.
Data export
@@
Data analysis
Once verified, these retrieved observations form the foundation for a wide range of statistical analyses, demonstrated in the following section. This workflow allows experts to effectively “design” their search strategy and specify what ecological signals to target, while identifying or mitigating relevant sources of uncertainty and bias.
Appendix B: Retrieval and filtering results
@@
American Robin “<species> with <diet type> in its mouth” Summer 500 per type 338 67.6%
@@
Young coniferous trees “young coniferous trees in burned forest” Geo: High Park Fire (40.57-40.75 N, 105.18-105.54 W); Date: Post-2012 100 45 45.0%
Young deciduous trees “young deciduous trees in burned forest” 112 78 69.6%
@@
Boston (Urban) “dead bird” Geo: Bounding box around Boston, MA; No Taxon filter 1000 543 54.3%
Pioneer Valley (Rural) “dead bird” Geo: Bounding box around Pioneer Valley/Amherst, MA; No Taxon filter 360 79 21.9%
@@
Emergence “Milkweed germinating or emerging” Taxon: Asclepias syriaca; Geo: S. Quebec (45.03 to 46.54 N, -74.68 to -71.66 W); No temporal filter 200 45 22.5%
Flowering “Milkweed flowering” 200 169 84.5%
Seeding “Milkweed producing seeds…” 200 161 80.5%
Senescence “Milkweed dying or withering…” 200 52 26.0%
@@
Humpback whale flukes “white underside of humpback whale fluke” Taxon: Humpback Whale (Megaptera novaeangliae); No Geo/Temp filters 200 153 76.5%
Humpback whale flukes “white underside of humpback whale fluke” Taxon: Humpback Whale (Megaptera novaeangliae); No Geo/Temp filters 200 153 76.5%